# LICO: Large Language Models for In-Context Molecular Optimization

### v1

@@ Abstract
Optimizing black-box functions is a fundamental problem in science and engineering. To solve this problem, many approaches learn a surrogate function that estimates the underlying objective from limited historical evaluations. Large Language Models (LLMs), with their strong pattern-matching capabilities via pretraining on vast amounts of data, stand out as a potential candidate for surrogate modeling. However, directly prompting a pretrained language model to produce predictions is not feasible in many scientific domains due to the scarcity of domain-specific data in the pretraining corpora and the challenges of articulating complex problems in natural language. In this work, we introduce LICO, a general-purpose model that extends arbitrary base LLMs for black-box optimization, with a particular application to the molecular domain. To achieve this, we equip the language model with a separate embedding layer and prediction layer, and train the model to perform in-context predictions on a diverse set of functions defined over the domain. Once trained, LICO can generalize to unseen molecule properties simply via in-context prompting. LICO achieves state-of-the-art performance on PMO, a challenging molecular optimization benchmark comprising over $20$ objective functions.
@@ 1 Introduction
Black-box optimization (BBO) is the problem of optimizing an unknown, often complex objective function without direct access to its structure or derivatives. This problem is ubiquitous in many science and engineering fields, including material discovery [18], protein engineering [8, 40, 2], molecular design [16], mechanical design [5, 30], and neural architecture search [52]. Typically, BBO involves an iterative process where each step constructs a surrogate model to approximate the objective function. This model then guides the selection of promising candidates for subsequent evaluation. The main challenge of this approach lies in learning an effective surrogate function that can accurately estimate the objective from limited historical data.
In stark contrast, we have seen impressive generalization abilities of Large Language Models (LLMs) [10, 1, 11, 43, 44, 45, 22, 23] for language-driven reasoning over many kinds of domains. By pretraining on Internet-scale data, LLMs have demonstrated exceptional pattern-matching abilities and generalization from limited observations in both natural language [10, 25, 48] and other domains [32, 35, 17]. This positions LLMs as a promising solution for enhancing surrogate modeling for BBO. Some recent works have indeed shown great potential for using LLMs for solving optimization problems [50, 12, 51, 3]. The main idea behind these methods is to frame the optimization problem in natural language, and prompt the language model using previously collected observations to make predictions for new data points [3] or to propose better candidates [50, 12, 51, 33, 38, 34, 28, 7, 31]. However, this approach has several limitations. First, performing optimization in the text space requires the problem and solution to be expressed in natural language, thus limiting this approach to selected domains. Second, the scarcity of domain-relevant data in the text corpora used to train language models poses generalization challenges when using these models for general scientific domains such as molecular optimization. Therefore, existing works have only demonstrated the success of LLMs in neural architecture search [3, 12, 51], prompt optimization [50], and code generation [33, 28], corresponding to domains that are well-represented in the training dataset for common language models [10, 44, 22]. Third, relying on verbose textual descriptions for both the problem and its solution imposes practical constraints by inflating the context length and thereby reducing the number of historical observations the model can effectively utilize.
@@
After training, LICO is capable of optimizing a wide range of molecular properties purely via in-context prompting. While the methodology of LICO applies to general scientific domains, in this paper we focus on molecular optimization. This problem plays a pivotal role in advancing drug and material discovery. The complexity of molecular structures and the vastness of the chemical space present unique challenges to black-box optimization algorithms. Moreover, since molecule-relevant data is likely under-represented in the pretraining corpora of existing language models, molecular optimization is a good problem to test the performance and applicability of LICO. We evaluate LICO against the state-of-the-art methods on Practical Molecular Optimization (PMO) [14], a challenging molecular optimization benchmark with over $20$ objective functions. The experiments show that LICO achieves the best performance and is the highest-ranked method in the benchmark.
@@ 2 Problem Statement
where we do not have direct access to the structure or gradient information of $f$. In molecular optimization, $\mathcal{X}$ is the space of all possible molecules, and $f$ is a certain property of the molecule we want to optimize over, such as bioactivities against a disease. While $f$ is unknown, we often have access to an unlabeled dataset $\mathcal{D}_{\text{u}}$ that consists of molecules $x^{\prime}s$ without the corresponding function values $y^{\prime}s$. ZINC [42] is such a dataset, which contains thousands to millions of unlabeled molecules.
To solve the optimization in (1), we can query $f$ with a limited budget, since evaluation often involves expensive physical experiments. To overcome this challenge, a common BBO approach learns a surrogate model $f_{\theta}$ that approximates the objective $f$ from past observations $\mathcal{D}_{\text{obs}}=\{(x_{i},y_{i})\}_{i=1}^{n}$, which starts empty and incrementally expands with new data points $(x,f(x))$ we query at each iteration. Formally, a surrogate model represents a predictive distribution $p_{\theta}(y\mid x,\mathcal{D}_{\text{obs}})$ of the function value $y$ conditioned on the input $x$ and the evolving observed dataset $\mathcal{D}_{\text{obs}}$. The prediction of this surrogate guides the selection of candidates to balance exploration and exploitation for subsequent function evaluation. The newly selected points are then added to $\mathcal{D}_{\text{obs}}$, and the process continues.
The success of this approach highly depends on the efficiency of the surrogate model $f_{\theta}$ in estimating $f$ from limited data in $\mathcal{D}_{\text{obs}}$ at each iteration. This resembles few-shot prediction, a setting that Large Language Models (LLMs) have proven to excel in. By pretraining on vast Internet-scale data, LLMs can learn generalizable patterns from limited data, and are capable of adapting to multiple functions at test time simply via in-context prompting [10, 35, 26, 27]. A recent line of works [50, 51, 12, 3] has exploited this ability of LLMs for optimization, but they relied on natural language as the interface, thus lacking generality to scientific domains. In this work, we propose a more general and efficient approach to leveraging LLMs for black-box optimization.
@@ 3 Related Work
LLMs for Optimization Recent works have explored the use of LLMs for optimization. The general idea behind these works is to prompt the model with the textual description of the optimization problem and historical evaluations for few-shot reasoning. Yang et al. [50], Liu et al. [31], Zhang et al. [51], Ma et al. [33] propose to prompt the language model to directly suggest better candidates to evaluate given the past inputs and their corresponding scores. Meyerson et al. [34], Lehman et al. [28], Bradley et al. [7] integrate LLMs with evolutionary algorithms, and prompt the model to perform crossover and mutation operations based on the population at each optimization step. Anonymous [3] study the use of LLMs to enhance several components in Bayesian optimization, including warmstarting, surrogate modeling, and candidate generation.
@@
The common approach in existing works has several inherent limitations. First, for general scientific domains, the input $x$ may not be easily described by natural language. Second, even when there is a textual description of the input, for instance, molecules can be represented by SMILES strings [49], it is likely that molecule-relevant data is significantly under-represented in the text corpora used for training existing language models. This hinders the generalization of LLMs to new domains outside of their training distribution. These challenges are why existing works only consider hyperparameter optimization, prompt optimization, and code generation, domains that are easy to describe in text, and are well-represented in the training data of LLMs. Furthermore, from an engineering perspective, naively prompting a language model with verbose textual descriptions of the input $x$ results in an excessively long context, thus reducing the number of examples the model can condition on. For example, an LLM with a maximum context length of $4000$ can only utilize up to $100$ past observations, assuming the average length of each data point is $40$. This practically limits the scalability of this approach to harder problems that require more steps to solve.
@@
LLMs for Non-language Tasks In addition to optimization, several works have studied the extension of pretrained LLMs to non-language domains with two main directions. The first direction considers problems that can be described in natural language, and prompts a pretrained LLM to solve the problem directly in the text space [35, 13, 17]. The second direction tackles more general problems, and does it by learning separate encoders for the new domain and aligning it with the embedding space of the pretrained LLM [32, 41, 47, 29]. Our work is closely related to the latter direction. However, as discussed in the following sections, while many of these works completely leave the word space, we find it beneficial to include language instruction while training the new modules.
@@ 4.1 Model Architecture
It is worth noting that the combination of natural language and domain-specific embeddings is the main distinction between LICO and previous works such as FPT [32] which applies pretrained LLMs to sequence classification tasks in non-language modalities. FPT also learns new embedding layers for the new domain, but relies entirely on the pretrained self-attention layers to model these embeddings without any language instructions. This distinction stems from the different nature of the tasks we aim to tackle. In sequence classification, the model produces a single prediction for the entire sequence, thus having a good representation of the sequence via self-attention is sufficient. For in-context learning, however, the model needs to associate each input $x$ with its corresponding value $y$ to infer the underlying function $f$ and make predictions for unknown $y$. A language instruction that specifies where $x$ is and where $y$ is helps the model identify this association and improve its in-context reasoning. Our ablation study in 5.2.1 confirms this utility of retaining language tokens.
@@ 4.2 Semi-synthetic Training
Our goal is to train LICO on the unlabeled data $\mathcal{D}_{\text{u}}$ with an unsupervised objective to facilitate efficient generalization to an arbitrary objective function $f$ in the same domain $\mathcal{X}$ after training. Our key insight is, that if we train the model to perform the estimation in (2) for a wide range of functions, it should be able to adapt to any objective function after training. While the true function values are unknown before optimization, we can use the unlabeled data $x^{\prime}s$ to generate training data from other functions. Assume we have access to a family of functions $\tilde{\mathcal{F}}$ that operate on the same input domain $\mathcal{X}$. For each function $\tilde{f}$ drawn from $\tilde{\mathcal{F}}$, we sample a set of function evaluations $\{(x_{i},y_{i})\}_{i=1}^{n}$ and train the model to autoregressively predict $y$ given the input $x$ and preceding $(x,y)$ pairs:
@@
in which the expectation is with respect to $\tilde{f}\sim\tilde{\mathcal{F}}$, $x_{1:n}\sim\mathcal{D}_{\text{u}}$, and $y_{1:n}=\tilde{f}(x_{1:n})$. After training, the estimation in (2) can be done purely via in-context prompting, where we condition the model on past observations to make predictions for new data points.
@@
$\tilde{f}\sim\mathcal{G}\mathcal{P}(0,\mathcal{K}),\hskip 5.69046pt\mathcal{K}(x,x^{\prime})=\frac{x\cdot x^{\prime}}{||x||^{2}+||x^{\prime}||^{2}-x\cdot x^{\prime}},$ (4)
@@
The final family of functions $\tilde{\mathcal{F}}$ used to train LICO is a mixture of intrinsic and synthetic functions with a certain ratio. This design choice is critical to the model’s performance. Intuitively, training on both types of functions ensures proximity to the downstream objectives and good coverage of the function space for efficient generalization. The use of intrinsic functions is also the main difference between our work and ExPT [37], a recent method that studies pure synthetic pretraining for optimization. We hypothesize that while synthetic data alone is sufficient for ExPT on a few simple tasks, for a more complex domain such as molecular optimization, synthetic training provides too little relevant signal for the model to generalize to downstream objectives. We empirically show the importance of both intrinsic and synthetic functions in the ablation study in section 5.2.2.
@@ 4.3 LICO for Black-box Optimization
After training, a single LICO model can be used for optimizing various objective functions within the domain $\mathcal{X}$. Optimization involves an iterative process. At each iteration $t$, we generate a set of candidates $\{x_{i}\}_{i=1}^{C}$ for which the model predicts the mean $\mu_{i}$ and standard deviation $\sigma_{i}$ conditioned on prior observations $\mathcal{D}_{\text{obs}}$, a dataset of $(x,y)$ pairs collected until $t$. An acquisition function $\alpha$ then calculates a utility score based on $\mu_{i}$ and $\sigma_{i}$ for each candidate, balancing between exploration (favoring high $\sigma$) and exploitation (favoring high $\mu$). The top $k$ candidates determined by their utility scores are evaluated using the objective function $f$. These $k$ candidates and their corresponding evaluations are incorporated into the dataset $\mathcal{D}_{\text{obs}}$, and the cycle repeats. This process terminates once we exhaust the evaluation budget of $B$. Algorithm 1 summarizes the optimization process.

### v2

@@ Abstract
Optimizing black-box functions is a fundamental problem in science and engineering. To solve this problem, many approaches learn a surrogate function that estimates the underlying objective from limited historical evaluations. Large Language Models (LLMs), with their strong pattern-matching capabilities via pretraining on vast amounts of data, stand out as a potential candidate for surrogate modeling. However, directly prompting a pretrained language model to produce predictions is not feasible in many scientific domains due to the scarcity of domain-specific data in the pretraining corpora and the challenges of articulating complex problems in natural language. In this work, we introduce LICO, a general-purpose model that extends arbitrary base LLMs for black-box optimization, with a particular application to the molecular domain. To achieve this, we equip the language model with a separate embedding layer and prediction layer, and train the model to perform in-context predictions on a diverse set of functions defined over the domain. Once trained, LICO can generalize to unseen molecule properties simply via in-context prompting. LICO performs competitively on PMO, a challenging molecular optimization benchmark comprising $23$ objective functions, and achieves state-of-the-art performance on its low-budget version PMO-1K.
@@ 1 Introduction
Black-box optimization (BBO) is the problem of optimizing an unknown, often complex objective function without direct access to its structure or derivatives. This problem is ubiquitous in many science and engineering fields, including material discovery (Hamidieh, 2018), protein engineering (Brookes et al., 2019; Sarkisyan et al., 2016; Angermueller et al., 2020), molecular design (Gaulton et al., 2012), mechanical design (Berkenkamp et al., 2016; Liao et al., 2019), and neural architecture search (Zoph & Le, 2016). Typically, BBO involves an iterative process where each step constructs a surrogate model to approximate the objective function. This model then guides the selection of promising candidates for subsequent evaluation. The main challenge of this approach lies in learning an effective surrogate function that can accurately estimate the objective from limited historical data.
In stark contrast, we have seen impressive generalization abilities of Large Language Models (LLMs) (Brown et al., 2020; Achiam et al., 2023; Bubeck et al., 2023; Team et al., 2023; Touvron et al., 2023a; b; Jiang et al., 2023; 2024) for language-driven reasoning over many kinds of domains. By pretraining on Internet-scale data, LLMs have demonstrated exceptional pattern-matching abilities and generalization from limited observations in both natural language (Brown et al., 2020; Kojima et al., 2022; Wei et al., 2022) and other domains (Lu et al., 2022; Mirchandani et al., 2023; Gruver et al., 2023). This positions LLMs as a promising solution for enhancing surrogate modeling for BBO. Some recent works have indeed shown great potential for using LLMs for solving optimization problems (Yang et al., 2023; Chen et al., 2023; Zhang et al., 2023; Liu et al., ). The main idea behind these methods is to frame the optimization problem in natural language, and prompt the language model using previously collected observations to make predictions for new data points (Liu et al., ) or to propose better candidates (Yang et al., 2023; Chen et al., 2023; Zhang et al., 2023; Ma et al., 2023; Nie et al., 2023; Meyerson et al., 2023; Lehman et al., 2023; Bradley et al., 2024; Liu et al., 2023a). However, this approach has several limitations. First, performing optimization in the text space requires the problem and solution to be expressed in natural language, thus limiting this approach to selected domains. Second, the scarcity of domain-relevant data in the text corpora used to train language models poses generalization challenges when using these models for general scientific domains such as molecular optimization. Therefore, existing works have only demonstrated the success of LLMs in neural architecture search (Liu et al., ; Chen et al., 2023; Zhang et al., 2023), prompt optimization (Yang et al., 2023), and code generation (Ma et al., 2023; Lehman et al., 2023), corresponding to domains that are well-represented in the training dataset for common language models (Brown et al., 2020; Touvron et al., 2023a; Jiang et al., 2023). Third, relying on verbose textual descriptions for the problem and its solution imposes practical constraints by inflating the context length and reducing the number of historical observations the model can effectively utilize.
@@
After training, LICO is capable of optimizing a wide range of molecular properties purely via in-context prompting. While the methodology of LICO applies to general scientific domains, in this paper we focus on molecular optimization. This problem plays a pivotal role in advancing drug and material discovery. The complexity of molecular structures and the vastness of the chemical space present unique challenges to black-box optimization algorithms. Moreover, since molecule-relevant data is likely under-represented in the pretraining corpora of existing language models, molecular optimization is a good problem to test the performance and applicability of LICO. We demonstrate the competitive performance of LICO against the leading methods on Practical Molecular Optimization (PMO) (Gao et al., 2022), a challenging molecular optimization benchmark with $23$ objective functions. On PMO’s low-budget setting, which we term PMO-1M, LICO achieves the best performance and is the highest-ranked method in the benchmark.
@@ 2 Problem Statement
where we do not have direct access to the structure or gradient information of $f$. In molecular optimization, $\mathcal{X}$ is the space of all possible molecules, and $f$ is a certain property of the molecule we want to optimize over, such as bioactivities against a disease. While $f$ is unknown, we often have access to an unlabeled dataset $\mathcal{D}_{\text{u}}$ that consists of molecules $x^{\prime}s$ without the corresponding function values $y^{\prime}s$. ZINC (Sterling & Irwin, 2015) is such a dataset with thousands of unlabeled molecules.
To solve the optimization in equation 1, we can query $f$ with a limited budget, since evaluation often involves expensive physical experiments. To overcome this challenge, a common BBO approach learns a surrogate model $f_{\theta}$ that approximates the objective $f$ from past observations $\mathcal{D}_{\text{obs}}=\{(x_{i},y_{i})\}_{i=1}^{n}$, which starts empty and incrementally expands with new data points $(x,f(x))$ we query at each iteration. Formally, a surrogate model represents a predictive distribution $p_{\theta}(y\mid x,\mathcal{D}_{\text{obs}})$ of the function value $y$ conditioned on the input $x$ and the evolving observed dataset $\mathcal{D}_{\text{obs}}$. The prediction of this surrogate guides the selection of candidates to balance exploration and exploitation. The newly selected points are added to $\mathcal{D}_{\text{obs}}$, and the process continues.
The success of this approach highly depends on the efficiency of the surrogate model $f_{\theta}$ in estimating $f$ from limited data in $\mathcal{D}_{\text{obs}}$ at each iteration. This resembles few-shot prediction, a setting that Large Language Models (LLMs) have proven to excel in. By pretraining on vast Internet-scale data, LLMs can learn generalizable patterns from limited data, and are capable of adapting to multiple functions at test time simply via in-context prompting (Brown et al., 2020; Mirchandani et al., 2023; Krishnamoorthy et al., 2023a; b). A recent line of works (Yang et al., 2023; Zhang et al., 2023; Chen et al., 2023; Liu et al., ) has exploited this ability of LLMs for optimization, but they relied on natural language as the interface, thus lacking generality to scientific domains. In this work, we propose a more general and efficient approach to leveraging LLMs for black-box optimization.
@@ 3 Related Work
LLMs for Optimization Recent works have explored the use of LLMs for optimization. The general idea behind these works is to prompt the model with the textual description of the optimization problem and historical evaluations for few-shot reasoning. Yang et al. (2023); Liu et al. (2023a); Zhang et al. (2023); Ma et al. (2023) propose to prompt the language model to directly suggest better candidates to evaluate given the past inputs and their corresponding scores. Meyerson et al. (2023); Lehman et al. (2023); Bradley et al. (2024) integrate LLMs with evolutionary algorithms, and prompt the model to perform crossover and mutation operations based on the population at each optimization step. Liu et al. study the use of LLMs to enhance several components in Bayesian optimization, including warmstarting, surrogate modeling, and candidate generation. Optformer (Chen et al., 2022) proposed to train an LLM specialized for in-context function prediction and optimization.
LLMs for Molecular Optimization Recent works have proposed to leverage LLMs for molecular optimization via prompting (Wang et al., 2024; Liu et al., 2023b; Ramos et al., 2023; Völker et al., 2024), leveraging LLM embeddings (Ranković & Schwaller, 2023), or finetuning on molecular corpora (Guevorguian et al., 2024; Ye et al., 2023; Fang et al., 2023; Kristiadi et al., ). MOLLEO (Wang et al., 2024) and ChatDrug (Liu et al., 2023b) are two prominent works in the first direction. MOLLEO proposed to prompt a pretrained LLM to perform crossover and mutation operations in a standard graph genetic algorithm, but its performance largely depends on the prompt format. Similarly, ChatDrug prompts a pretrained LLM for drug editing and requires a retrieval database of molecules to inject domain feedback into the LLM. BoChemian (Ranković & Schwaller, 2023) and Kristiadi et al. studied the use of LLM embeddings for Bayesian molecular optimization, and Kristiadi et al. additionally explored finetuning an LLM to serve as a surrogate for optimization. Chemlactica/Chemma (Guevorguian et al., 2024) and MOLGEN (Fang et al., 2023) proposed to pretrain and/or finetune LLMs on molecule-related corpora to generate valid molecules, which can serve as a genetic algorithm in molecular optimization. DrugAssist (Ye et al., 2023) creates the MolOpt-Instructions dataset that contains pairs of molecules and their property values to finetune a pretrained LLM that can iteratively propose better molecules after training.
@@
The common approach in existing works has several inherent limitations. First, for general scientific domains, the input $x$ may not be easily described by natural language. Second, even when there is a textual description of the input, for instance, molecules can be represented by SMILES strings (Weininger, 1988), existing prompt-based works require significant prompt optimization to achieve good performance, and the optimal prompt often varies between tasks. Furthermore, from an engineering perspective, naively prompting a language model with verbose textual descriptions of the input $x$ results in an excessively long context, thus reducing the number of examples the model can condition on. For example, an LLM with a maximum context length of $4000$ can only utilize up to $100$ past observations, assuming the average length of each data point is $40$. This practically limits the scalability of this approach to harder problems that require more steps to solve.
@@
LLMs for Non-language Tasks In addition to optimization, several works have studied the extension of pretrained LLMs to non-language domains with two main directions. The first direction considers problems that can be described in natural language, and prompts a pretrained LLM to solve the problem directly in the text space (Mirchandani et al., 2023; Dinh et al., 2022; Gruver et al., 2023; Liu et al., 2024; Sprueill et al., 2024). The second direction tackles more general problems by learning separate encoders for the new domain and aligning it with the embedding space of the pretrained LLM (Lu et al., 2022; Shen et al., 2023; Tsimpoukelli et al., 2021; Li et al., 2022). Our work is closely related to the latter direction. However, as discussed in the following sections, while many of these works completely leave the word space, we find it beneficial to include language instruction while training the new modules.
@@ 4.1 Model Architecture
It is worth noting that the combination of natural language and domain-specific embeddings is the main distinction between LICO and previous works such as FPT (Lu et al., 2022) which applies pretrained LLMs to sequence classification tasks in non-language modalities. FPT also learns new embedding layers for the new domain, but relies entirely on the pretrained self-attention layers to model these embeddings without any language instructions. This distinction stems from the different nature of the tasks we aim to tackle. In sequence classification, the model produces a single prediction for the entire sequence, thus having a good representation of the sequence via self-attention is sufficient. For in-context learning, however, the model must associate each input $x$ with its value $y$ to infer the underlying function $f$ and make predictions for unknown $y$. A language instruction that specifies where $x$ is and where $y$ is helps the model identify this association and improve its in-context reasoning. Our ablation study in 5.2.1 confirms this utility of retaining language tokens.
@@ 4.2 Semi-synthetic Training
Our goal is to train LICO on the unlabeled data $\mathcal{D}_{\text{u}}$ with an unsupervised objective to facilitate efficient generalization to an arbitrary objective function $f$ in the same domain $\mathcal{X}$ after training. Our key insight is that if we train the model to perform the estimation in equation 2 for a wide range of functions, it should adapt to any objective function post-training. While the true function values are unknown before optimization, we can use the unlabeled data $x^{\prime}s$ to generate training data from other functions. Assume we have access to a family of functions $\tilde{\mathcal{F}}$ that operate on the same input domain $\mathcal{X}$. For each function $\tilde{f}$ drawn from $\tilde{\mathcal{F}}$, we sample a set of function evaluations $\{(x_{i},y_{i})\}_{i=1}^{n}$ and train the model to autoregressively predict $y$ given the input $x$ and preceding $(x,y)$ pairs:
@@
in which the expectation is with respect to $\tilde{f}\sim\tilde{\mathcal{F}}$, $x_{1:n}\sim\mathcal{D}_{\text{u}}$, and $y_{1:n}=\tilde{f}(x_{1:n})$. After training, the estimation in equation 2 can be done purely via in-context prompting, where we condition the model on past observations to make predictions for new data points.
@@
[…]
@@
The final family of functions $\tilde{\mathcal{F}}$ used to train LICO is a mixture of intrinsic and synthetic functions with a certain ratio. This design choice is critical to the model’s performance. Intuitively, training on both types of functions ensures proximity to the downstream objectives and good coverage of the function space for efficient generalization. The use of intrinsic functions is also the main difference between our work and ExPT (Nguyen et al., 2023), a recent method that studies pure synthetic pretraining for optimization. We hypothesize that while synthetic data is sufficient for ExPT on a few simple tasks, for a more complex domain such as molecular optimization, synthetic training provides too little relevant signal for the model to generalize to downstream objectives. We empirically show the importance of both intrinsic and synthetic functions in the ablation study in section 5.2.2.
@@ 4.3 LICO for Black-box Optimization
After training, a single LICO model can be used for optimizing various objective functions within the domain $\mathcal{X}$. Optimization involves an iterative process. At each iteration $t$, we generate a set of candidates $\{x_{i}\}_{i=1}^{C}$ using standard crossover and mutation operations for which the model predicts the mean $\mu_{i}$ and standard deviation $\sigma_{i}$ conditioned on prior observations $\mathcal{D}_{\text{obs}}$, a dataset of $(x,y)$ pairs collected until $t$. An acquisition function $\alpha$ then calculates a utility score based on $\mu_{i}$ and $\sigma_{i}$ for each candidate, balancing between exploration (favoring high $\sigma$) and exploitation (favoring high $\mu$). The top $k$ candidates determined by their utility scores are evaluated using the objective function $f$. These $k$ candidates and their corresponding evaluations are incorporated into the dataset $\mathcal{D}_{\text{obs}}$, and the cycle repeats. This process terminates once we exhaust the evaluation budget of $B$. Algorithm 1 summarizes the optimization process and Appendix A.3 outlines the optimization hyperparameters.